# Mlynská dolina, 842 48, Bratislava, Slovakia

### v1

@@ Abstract
The fuzzy onion model formed by connecting a set of concentric fuzzy spheres of increasing radius is motivated by studies of quantum space but can also be used to study standard physics. The main feature of the model is that functions in three-dimensional space, for example, scalar fields or wavefunctions, are expressed in terms of Hermitian matrices of a certain form. Relevant equations are then matrix equations, and some problems, such as searching for energy spectrum for fixed angular momentum quantum numbers $(l,m)$, can be expressed as an eigenvalue problem. We show how this simple approach can reproduce the results of other studies analyzing the hydrogen atom in a spherical cavity. We also test the effect of short distance quantum structure of the space on these solutions — not looking for the phenomenological consequences, as the scale of quantum space is many orders below the order of the Bohr radius, but to understand the effect quantum space in general. We observe a set of solutions with no classical counterpart that has been described in a former theoretical study.
@@ 1 Introduction
The hydrogen atom is one of the most important objects in physics. This is not only due to its abundance, but it has been thoroughly studied experimentally, analytically, and numerically. In quantum mechanics, it serves as important role as black holes do in general relativity — it is a rare real example of an exactly solvable problem. In this paper, we present a solution to the problem of hydrogen atoms confined to a spherical cavity. We came across this solution when analysing the behaviour of quantum space, a mathematical formulation that seeks to describe the physics of the Planck length scale.
The Planck scale is obtained by combining the fundamental physical constants, $G,\hbar,c$, into new ones with the dimensions of length, mass, energy and so on. From their values, one can infer that the space should not be continuous as is usually assumed in physical models but obtains a quantum structure on the scale of $\mbox{l}_{\scriptsize{Pl}}\approx 10^{-35}m$. Models of quantum space have been considered in the earliest days of quantum mechanics, perhaps as a regularisation tool, but were eclipsed by the novel methods of renormalization. They were brought back to the spotlight when the community started discussing models of quantum gravity, for which scales given by all three aforementioned constants are relevant and appear in certain aspects of the string theory and matrix models of M-theory. This theory has not yet been constructed; there are various promising ideas and approaches doplicher ; Chamseddine:1996zu ; Ambjorn:2004qm ; Szabo:2001kg ; Steinacker:2011ix ; Seiberg:1999vs ; Berenstein:2015pxa ; hsjt22 and many of them share a common feature of quantum space Hossenfelder:2012jw ; Nicolini:2022rlz .
A quantum space is not a completely well-defined term, as different lines of research perceive it differently. A common way of thinking about it is that it is a space with limited spatial resolution. Any experiment trying to distinguish points below a certain scale, defined by a parameter of length here denoted $\lambda$, which is reasonably assumed to be of the Planck length scale, will fail Hossenfelder:2012jw . For example, a particle with a wavelength short enough to distinguish between such points would form a black hole, hiding any desired information.
@@
An important model of quantum space is the fuzzy sphere, which is obtained by putting an upper cut-off on the spectrum of angular momenta and describing the fields using Hermitian matrices. The size of the matrix is inversely proportional to the spatial resolution — the larger the matrix size, the higher angular momenta are allowed, and one is able to approximate a positional delta function more and more accurately. The fuzzy sphere model has been studied from various points of view Hoppe ; Madore:1991bw since it is among the simplest and most instructive models of quantum space, despite the fact that some constructional issues remain unresolved diffstar ; Kovacik:2018thy ; stripes .
Recently, we have formulated a three-dimensional space that is formed as a sequence of concentric fuzzy spheres of increasing radius Kovacik:2023zab , built along lines similar to other works seik ; pgk15 ; pg13 ; todorovic1 ; todorovic2 ; sch1 ; sch2 ; sch3 ; sch4 ; patricia ; patricia2 ; Tekel:2015uza . The main goal was not to obtain a precise mathematical description of quantum space originating from a fundamental theory but to obtain a model that carries both the essential features of quantum space and is reasonably amendable for studies of interesting problems — field theories, classical mechanics, waves on neutron stars, gravitational collapses and models of microscopic black holes. This wide range of areas of applicability stems from the fact that one can change the value of $\lambda$ to a different value, and then it, instead of being the Planck scale, can describe a scale of the granularity of the given material granular .
@@ 2 A brief overview of the fuzzy onion model
where differential operators act via the commutator, $\hat{L}_{i}Y=[L_{i},Y]$. These form an infinite-dimensional representation of the $su(2)$ algebra. There are, however, finite-dimensional representations; the smallest nontrivial one expressed in terms of Pauli matrices might be the most known one. For any matrix size $N$, one can find three matrices $L_{i}$ and $N^{2}$ matrices $Y_{lm}$ that satisfy relations (1); details can be found in the appendix A. The matrices $Y_{lm}$ form a basis for Hermitian matrices of the given size, which means any Hermitian matrix can be expanded as
@@
where $l=0,1,\ldots,N-1$ and $m=-l,\ldots,l$. This means that such matrices can be mapped onto functions of angular coordinates $\Phi\rightarrow\phi(\theta,\varphi)$ just by using the same expansion coefficients. The map works both ways as long as we use only functions with maximal momentum in the expansion equal to $N-1$; such functions cannot have arbitrarily small support — this is what is meant by the limited spatial resolution. Such functions can be mapped to Hermitian matrices that form a noncommutative algebra closed under multiplication with a defined set of rotation generators. Other relevant objects can be defined as well; for example, taking the integral of $\phi$ corresponds to taking the trace of $\Phi$. A lot of physics can be studied on the fuzzy sphere; we refer the reader to Kovacik:2023zab and move to the construction of the fuzzy onion model.
The angular momenta in the expression (1) are dimensionless, a noncommutative algebra defined by $[x_{i},x_{i}]=2i\lambda\varepsilon_{ijk}x_{k}$ requires a new parameter with the dimension of length, which is denoted $\lambda$ and defined as $x_{i}=\lambda L_{i}$. From this, one can derive a relation between the matrix size and radius of the fuzzy sphere: $R=\lambda\sqrt{N^{2}-1}\approx\lambda N$; here, we take $R=\lambda N$. That means that by keeping $\lambda$ fixed, larger matrices correspond to larger spheres. That is, a sphere of radius $\lambda$ can be described by a matrix of size one, a sphere of radius $2\lambda$ by a $2\times 2$ matrix and so on. That means that the set of matrices of increasing size, $\left(\Phi^{(1)},\Phi^{(2)},\ldots,\Phi^{(M)}\right)$, where the superscript denotes the matrix size, describe a set of $M$ disjoint fuzzy spheres, which can be taken to be concentric at this point. This set can be placed into a larger matrix $\Psi$ defined as
@@
[…]
$\int d^{3}\psi\leftrightarrow Tr\left(4\pi r\Psi\right),$ (4)
where $r$ is a diagonal matrix with entries $\lambda,2\lambda,2\lambda,3\lambda,3\lambda,3\lambda,4\lambda\ldots$ As we can define the angular part of the Laplace operator, that is $\Delta_{L}=-\frac{\hat{{\cal{L}}}}{r^{2}}$ on each of the spheres, it can be defined on the matrix $\Psi$ as well:
${\cal{L}}\Psi=\begin{pmatrix}\hat{{\cal{L}}}^{(1)}\Phi^{(1)}&&&\\ &\hat{{\cal{L}}}^{(2)}\Phi^{(2)}&&\\ &&\ddots&\\ &&&\hat{{\cal{L}}}^{(M)}\Phi^{(M)}.\end{pmatrix}.$ (5)
Defining the redial part of the kinetic term requires a certain nuance, as one cannot directly subtract matrices of different sizes as is needed when computing the radial derivative. This issue can be solved by defining two operators, ${\cal{U}}$ and ${\cal{D}}$ that decrease and increase the matrix size; details can be found in the appendix B. Using those, we can define
@@
Using those, the full Laplace operator acting on the fuzzy onion states can be defined as
${\cal{K}}\Psi=\left(\frac{1}{r^{2}}\frac{\partial}{\partial_{r}}\left(r^{2}\frac{\partial}{\partial r}\right)-\frac{{\cal{L}}^{2}}{r^{2}}\right)\Psi,$ (8)
which serves as the kinetic energy operator in the Hamiltonian. We are now ready to analyze quantum mechanical problems in this space.
@@ 3.1 Construction and standard states
where $V(r)=0$ for $r<r_{0}$ and ${V(r)=\infty}$ otherwise, here $r_{0}$ is the size of a spherical cavity. Here, $q$ and $m$ are the charge and the mass of the electron, respectively, and will be set to $1$ from now on, thus setting the Bohr radius to $1$ as well; we also use units where $\hbar=1$ which also sets the Hartree energy to $1$.
In our construction, the cavity size is given by $r_{0}=M\lambda$, where $M$ is the size of the matrix describing the outermost layer. This can be viewed as either a fundamental quantum feature of the space or as a regularisation tool on our way to the $M\to\infty$ and $\lambda\to 0$ limit with fixed $r_{0}$ for a continuous space.
When solving this problem in standard quantum mechanics, one can exploit the rotational symmetry and split the solution into the form $\psi(r,\theta,\phi)=R(r)Y_{lm}(\theta,\phi)$. Something similar can be done in the case of the fuzzy onion model. When fixed values of $(l,m)$ are picked, then each layer is represented by a single coefficient $c^{(i)}_{lm}$ and they form a vector, dropping the trivial terms:
@@
we denote the operators acting on vectors $\mathcal{C}$ rather than matrices $\Phi$ in boldface and ${\cal{K}}_{R}$ is specified in the appendix B. Putting all the terms together, one obtains a single large matrix $\mathbf{H}$, its eigenvalues are the energy eigenvalues of the Hamiltonian, and the eigenvector forms the radial part of the solution.
We observe that eigenvalues generally split into two classes: one that corresponds to the bounded states and one that corresponds to scattering states, or to be more precise — the states that would correspond to those in the infinity-cavity limit. It is however not clear where the split between these two happens for a finite $M$. We have evaluated the eigenvalues for various combinations of $r_{0},M,n,l$ and observed a considerable agreement with the reference aquino2007highly ; the comparison is shown in tables in the section 3.3. The eigenvectors corresponding to the Hamiltonian eigenvalues are the radial parts of the energy eigenstates, and we can use mapping (2) to reconstruct the radial probability distribution for the electron. Two examples are shown in the figures 3 and 4.
We can also use them to compute the mean value of observables defined as
@@
Note that the observables are, as other relevant objects, matrices. As an example, we compute $\left\langle r\right\rangle$; the results are shown in tables 2 and 2. We show the results for two different values of $M$, which, when comparing the same values of $r_{0}$, relate to two different values of $\lambda$. It is clear and consistent that by increasing the value of $\lambda$, the particle is pushed outward from the centre. This suggests that the quantumness of space creates a form of repulsion.
Let us now consider the continuous space limit. If we keep $r_{0}$ fixed and increase $M$, the scale of noncommutative $\lambda$ vanishes, and we reproduce the hydrogen atom in the spherical cavity of radius $r_{0}$ in ordinary quantum mechanics. One can also do a double limit and send $M\rightarrow\infty$, which should reproduce the ordinary hydrogen atom spectrum. When studying the confined hydrogen, three regimes are of interest: a strongly compressed particle with $r_{0}<a_{0}$, where $a_{0}$ is the Bohr radius, equal to unity in our units. The second regime is a mildly compressed atom with $r_{0}\approx a_{0}$ and then a barely confined atom with $r_{0}\gg a_{0}$. These, all investigated in the commutative limit, are interesting from a practical point of view in a chemical context Reyes2024 ; Scherbinin1998 ; AlHashimi2015 ; Reyes2024a ; AlHashimi2012 ; Kamel2016 ; Goldman1992 ; Aquino2009 ; Laughlin2002 ; White2015 ; Ping2019 ; Chaudhuri2021 ; Fischer2023 ; aquino2007highly .
3.2 Spurious states
We are also interested in the regime of strong effects of quantum space. We can again consider three situations with regard to the size of confinement, but now, instead of taking $\lambda\rightarrow 0$, we take $\lambda\rightarrow a_{0}$. Why is this situation interesting? Firstly, it can give some clues about the behaviour of particles in the earliest stages of the universe, where everything was very dense and squeezed and also energetic enough to make the effects of quantum space more prominent. Secondly, it can provide some insight into gravitational collapse. Of course, in our situation, there is no gravity, but earlier studies Nicolini:2005vd suggest that the effects of quantum space prevent the formation of a perfect singularity and provide a form of outward pressure that stabilizes the core so we can now see whether the quantumness of space would act against squeezing effort of the confinement.
We have also observed a class of solutions with no counterpart in the case of ordinary space. These have been conjectured in a previous study of the Hydrogen atom in a noncommutative space. It has been observed that in addition to states with energies $E_{n}^{I}$ that reproduce the spectrum of an unconfined Hydrogen atom in ordinary space, there is an additional set of bound states for a repulsive potential of the same form. The value of these energies is
$E_{n}^{II}=\frac{2}{\lambda^{2}}-E_{n}^{I}.$ (14)
A possible interpretation is that these states are a form of doublers resulting from the minimal-length scale. Another, perhaps related, interpretation is the duality discussed in Nicolini:2022rlz . In Table 5, we compare the two energy scales obtained as smallest and largest eigenvalues of 11 with the values predicted by 14 — the match is nearly perfect. In Figure 5, we compare the forms of solutions; again, any difference is difficult to observe. Therefore, we conclude the mirror form of the energy spectrum of 11 and the presence of spurious states. In Figure 2, we plot the radial part of the wavefunctions for an attractive potential. We obtained essentially identical results for repulsive potential, taking 14 into account.
An important aspect of our construction is that it can be easily modified to include various terms in the potential. To demonstrate this, we have added a linear term
$H_{\scriptsize{add}}=-5r,$ (15)
to the Hamiltonian (11) and found its spectrum and set of solutions, the first four of them are shown in figure 1. This potential has been analysed in the context of the study of bound states of quarks using different methods in Bukor:2022xjc .

### v2

@@ Abstract
The fuzzy onion model formed by connecting a set of concentric fuzzy spheres of increasing radius is motivated by studies of quantum space but can also be used to study standard physics. The main feature of the model is that functions in three-dimensional space – like scalar fields or wavefunctions – are expressed in terms of Hermitian matrices of a certain structure. Relevant equations are then matrix equations, and some problems, such as searching for the energy spectrum for fixed quantum numbers $(l,m)$, can be expressed as an eigenvalue problem. We show how this simple approach can reproduce the results of other studies analyzing the hydrogen atom in a spherical cavity. We also test the effect of the short-distance quantum structure of the space on these solutions — not looking for the phenomenological consequences, as the scale of quantum space is many orders below the order of the Bohr radius, but to understand the effect of quantum space in general. We observe a set of solutions without a classical counterpart which have been suggested also in a former theoretical study.
@@ 1 Introduction
The hydrogen atom is one of the most important objects in physics. Thanks to its abundance, it has been thoroughly studied experimentally, analytically, and numerically. In quantum mechanics, it serves an important role as black holes do in general relativity — it is a rare real example of an exactly solvable problem. In this paper, we present a solution to the problem of hydrogen atoms confined to a spherical cavity. We came across this solution when analysing the behaviour of quantum space, a mathematical formulation that seeks to describe the physics of the Planck length scale.
The Planck scale is obtained by combining the fundamental physical constants, $G,\hbar,c$, into objects with dimensions of length, mass, energy, etc. From their values, one can infer that the space should not be continuous as is usually assumed in physical models but obtains a quantum structure on the scale of $\mbox{l}_{\scriptsize{Pl}}\approx 10^{-35}m$. Models of quantum space have been considered in the earliest days of quantum mechanics, perhaps as a regularisation tool, but were eclipsed by methods of renormalization. They were brought back to the spotlight when the community started discussing models of quantum gravity, for which scales given by all three aforementioned constants are relevant. Moreover, they appear in certain aspects of the string theory and matrix models of M-theory. This theory has not yet been constructed; there are various promising ideas and approaches doplicher ; Chamseddine:1996zu ; Ambjorn:2004qm ; Szabo:2001kg ; Steinacker:2011ix ; Seiberg:1999vs ; Berenstein:2015pxa ; Steinacker:2022kji and many of them share a common feature of quantum space Hossenfelder:2012jw ; Nicolini:2022rlz . A quantum space is not a completely well-defined term, as different lines of research perceive it differently. A common way of thinking about it is that it is a space with limited spatial resolution. Any experiment trying to distinguish points below a certain scale, defined by a parameter of length here denoted $\lambda$, which is reasonably assumed to be of the Planck length scale, will fail Hossenfelder:2012jw . For example, a particle with a wavelength short enough to distinguish between such points would form a black hole, hiding any desired information.
@@
An important model of quantum space is the fuzzy sphere, which is obtained by putting a cut-off on the spectrum of angular momenta and describing the fields using Hermitian matrices. The size of the matrix is inversely proportional to the spatial resolution — the larger the matrix size, the higher angular momenta are allowed, and one is able to approximate a positional delta function more and more accurately. The fuzzy sphere model has been studied from various points of view Hoppe ; Madore:1991bw since it is among the simplest and most instructive models of quantum space, despite the fact that some constructional issues remain unresolved.
Recently, we have formulated a three-dimensional space that is formed as a sequence of concentric fuzzy spheres of increasing radius Kovacik:2023zab , built along lines similar to other works seik ; pgk15 ; pg13 ; Kupriyanov:2012nb ; todorovic1 ; todorovic2 ; sch1 ; sch2 ; sch3 ; sch4 ; patricia ; patricia2 ; Tekel:2015uza . The main goal was not to obtain a precise mathematical description of quantum space originating from a fundamental theory, but to obtain a model that carries both the essential features of quantum space and is reasonably amenable to studies of interesting problems — field theories, classical mechanics, waves on neutron stars, gravitational collapses and models of microscopic black holes. This wide range of areas of applicability stems from the fact that one can change the value of $\lambda$ to a different value, and then it can describe a scale of the granularity of the given material rather than the Planck scale granular .
@@ 2 A brief overview of the fuzzy onion model
where differential operators act via the commutator, $\hat{L}_{i}Y=[L_{i},Y]$. These form an infinite-dimensional representation of the $su(2)$ algebra. There are, however, finite-dimensional representations; the smallest nontrivial one expressed in terms of Pauli matrices might be the most well-known one. For any matrix size $N$, one can find three matrices $L_{i}$ and $N^{2}$ matrices $Y_{lm}$ that satisfy relations (1); details can be found in the appendix A. The matrices $Y_{lm}$ form a basis for Hermitian matrices of the given size, which means any Hermitian matrix can be expanded as
@@
where $l=0,1,\ldots,N-1$ and $m=-l,\ldots,l$. This means that such matrices can be mapped onto functions of angular coordinates $\Phi\rightarrow\phi(\theta,\varphi)$ just by using the same expansion coefficients. The map works both ways as long as we use only functions with maximal momentum in the expansion equal to N1N-1; such functions cannot have arbitrarily small support — this is what is meant by the limited spatial resolution. These functions can be mapped to Hermitian matrices that form a noncommutative algebra closed under multiplication with a defined set of rotation generators. Other relevant objects can be defined as well; for example, taking the integral of $\phi$ corresponds to taking the trace of $\Phi$. We refer the reader to Kovacik:2023zab for a more detailed discussion and move to a brief description of the fuzzy onion model.
The angular momenta in the expression (1) are dimensionless, a noncommutative algebra defined by $x_{i},x_{i}]=2i\lambda\varepsilon_{ijk}x_{k}$ requires a new parameter with the dimension of length, which is denoted $\lambda$ and defined as $x_{i}=2\lambda L_{i}$. From this, one can derive a relation between the matrix size and radius of the fuzzy sphere: $R=\lambda\sqrt{N^{2}-1}\approx\lambda N$; here, we take $R=\lambda N$. By keeping $\lambda$ fixed, larger matrices correspond to larger spheres. That is, a sphere of radius $\lambda$ can be described by a matrix of size one, a sphere of radius $2\lambda$ by a $2\times 2$ matrix and so on. That means that the set of matrices of increasing size, $\left(\Phi^{(1)},\Phi^{(2)},\ldots,\Phi^{(M)}\right)$, where the superscript denotes the matrix size, describe a set of $M$ disjoint concentric fuzzy spheres. This set can be placed into a larger matrix $\Psi$ defined as
@@
As the integration on a single fuzzy sphere is replaced by a trace, the same can be done here. There is an important detail — since $\mbox{Tr}(\ldots)$ corresponds to $\int(\ldots)r\ d\Omega$, one needs to add a certain factor to obtain the proper measure:
$\int d^{3}\psi\leftrightarrow\mbox{Tr}\left(4\pi r\Psi\right),$ (4)
where $r$ is a diagonal matrix with entries $\lambda,2\lambda,2\lambda,3\lambda,3\lambda,3\lambda,4\lambda,\ldots$ . As we can define the angular part of the Laplace operator, that is $\Delta_{L}=-\frac{{\cal{L}}}{r^{2}}$, on each of the spheres individually, it can be defined on the entire matrix $\Psi$ as well:
${\cal{L}}\Psi=\begin{pmatrix}\hat{{\cal{L}}}^{(1)}\Phi^{(1)}&&&\\ &\hat{{\cal{L}}}^{(2)}\Phi^{(2)}&&\\ &&\ddots&\\ &&&\hat{{\cal{L}}}^{(M)}\Phi^{(M)}\end{pmatrix}$. (5)
Defining the radial part of the kinetic term requires a certain nuance, as one cannot directly subtract matrices of different sizes, as is needed when computing the radial derivative. This issue can be solved by defining two operators, $\cal{U}$} and $\cal{D}$} that decrease and increase the matrix size; details can be found in the appendix B. Using those, we can define
@@
Then, the full Laplace operator acting on the fuzzy onion states can be defined as
${\cal{K}}\Psi=\left(\frac{1}{r^{2}}\frac{\partial}{\partial_{r}}\left(r^{2}\frac{\partial}{\partial r}\right)-\frac{{\cal{L}}}{r^{2}}\right)\Psi,$ (8)
which serves as the kinetic energy operator in the Hamiltonian. We are now ready to analyze the problem of the Hydrogen atom in the cavity in this setting.
@@ 3 The hydrogen atom in the spherical cavity
In this section, we describe our calculation and illustrate the most important findings. More detailed results are discussed in the appendix C.
@@ 3.1 Construction and standard states
where $V(r)=0$ for $r<r_{0}$ and $V(r)=\infty$} otherwise and $r_{0}$ is the size of a spherical cavity. Here, $q$ and $m$ are the charge and the mass of the electron, respectively, and will be set to $1$ from now on; we also use units where $\hbar=1$, which also sets the Bohr radius and the Hartree energy to $1$.
In our construction, the cavity size is given by $r_{0}=M\lambda$, where $M$ is the size of the matrix describing the outermost layer. This can be viewed as either a fundamental quantum feature of the space or as a regularisation tool for the continuous space limit obtained by taking $M\to\infty$ and $\lambda\to 0$ while keeping $r_{0}$ fixed.
When solving this problem in standard quantum mechanics, one can exploit the rotational symmetry and split the solution into the form $\psi(r,\theta,\phi)=R(r)Y(\theta,\phi)$. Something similar can be done in the case of the fuzzy onion model. When fixed values of $(l,m)$ are chosen, then each layer is represented by a single coefficient $c^{(i)}_{lm}$ and together they form a vector, dropping the trivial terms:
@@
we denote the operators acting on vectors $\mathcal{C}$ rather than matrices $\Phi$ in boldface and ${\cal{K}}_{R}$ is specified in the appendix B. Putting all the terms together, one obtains a single large matrix $\mathbf{H}$, its eigenvalues are the energy eigenvalues of the Hamiltonian, and the eigenvectors form the radial part of the solution.
We observe that eigenvalues generally split into two classes: one that corresponds to the bounded states and one that corresponds to scattering states, or to be more precise — the states that would correspond to those in the infinity-cavity limit; we separate them by the sign of their energy. We have evaluated the eigenvalues for various combinations of $r_{0},M,n,l$ and observed a considerable agreement with the reference aquino2007highly ; the comparison is shown in tables in the section 3.2. The eigenvectors corresponding to the Hamiltonian eigenvalues are the radial parts of the energy eigenstates, and we can use mapping (2) to reconstruct the radial probability distribution for the electron. Two examples are shown in Figures 1 and 6. We can see that our matrix description of the model can be used to efficiently reproduce the results from ordinary quantum mechanics.
We can compute the mean value of observables defined as
@@
Note that the observables are, like other relevant objects, matrices. As an example, we compute $\left\langle r\right\rangle$; the results are shown in tables 3 and 3. We show the results for two different values of $M$, which, when comparing the same values of $r_{0}$, relate to two different values of $\lambda$. It is clear and consistent that by increasing the value of $\lambda$, the particle is pushed outward from the centre. This suggests that the quantumness of space creates a form of repulsion. This insight can be useful for models of matter singularities in a quantum space.
Let us now consider the continuous space limit. If we keep $r_{0}$ fixed and increase $M$, the scale of noncommutative $\lambda$ vanishes, and we reproduce the model of the hydrogen atom in the spherical cavity of radius $r_{0}$ in ordinary quantum mechanics. One can also do a double limit and send $M\rightarrow\infty$, which reproduces the ordinary hydrogen atom spectrum. When studying the confined hydrogen, three regimes are of interest: a strongly compressed particle with $r_{0}<a_{0}$, where $a_{0}$ is the Bohr radius, equal to unity in our units. The second regime is a mildly compressed atom with $r_{0}\approx a_{0}$ and then a barely confined atom with $r_{0}\gg a_{0}$. These, all investigated in the commutative limit, are interesting from a practical point of view in a chemical context Reyes2024 ; Scherbinin1998 ; AlHashimi2015 ; Reyes2024a ; AlHashimi2012 ; Kamel2016 ; Goldman1992 ; Aquino2009 ; Laughlin2002 ; White2015 ; Ping2019 ; Chaudhuri2021 ; Fischer2023 ; aquino2007highly . Basically, the hydrogen atom placed in a cavity is a useful model to study changes in reactions confined to a small reaction site or under high pressure. Therefore, it might be relevant to analyze more complex chemical compounds in this setting; we leave this for future study.